# MC-NeRF: Multi-Camera Neural Radiance Fields for Multi-Camera Image Acquisition Systems

### v2

@@ 3.2.2   Training Parameters and Degenerated Cases
$A={{\left[\begin{matrix}\begin{matrix}\begin{matrix}-{{X}_{0}}\\ -{{Y}_{0}}\\ \end{matrix}\\ 0\\ 1\\ 0\\ \end{matrix}&\begin{matrix}\begin{matrix}0\\ 0\\ \end{matrix}\\ 0\\ 0\\ -{{X}_{0}}\\ \end{matrix}&\begin{matrix}\begin{matrix}\ldots\\ \ldots\\ \end{matrix}\\ \ldots\\ \ldots\\ \ldots\\ \end{matrix}\\ \begin{matrix}0\\ 0\\ \end{matrix}&\begin{matrix}-{{Y}_{0}}\\ 0\\ \end{matrix}&\begin{matrix}\ldots\\ \ldots\\ \end{matrix}\\ \begin{matrix}0\\ -{{u}_{0}}{{X}_{0}}\\ -{{u}_{0}}{{Y}_{0}}\\ \begin{matrix}0\\ {{u}_{0}}\\ \end{matrix}\\ \end{matrix}&\begin{matrix}1\\ {{v}_{0}}{{X}_{0}}\\ {{v}_{0}}{{Y}_{0}}\\ \begin{matrix}0\\ {{v}_{0}}\\ \end{matrix}\\ \end{matrix}&\begin{matrix}\ldots\\ \ldots\\ \ldots\\ \begin{matrix}\ldots\\ \ldots\\ \end{matrix}\\ \end{matrix}\\ \end{matrix}\right]}^{T}}$ (10)
In Eq.(10), columns 3,7 and 11 are always equal to 0, which implies that there is no unique solution for $A$ and also for $H$. This prevents available solution for the camera intrinsic matrix ${K_{cam}^{i}}$. The specific degenerated case and normal case are shown in Fig.3.
3.2.3   Calibration cube design and instruction
Compared to single-camera calibration, multi-camera calibration has the following limitations: First, the cameras of multi-camera systems are fixed in space and cannot move freely, like camera array and camera ball. Second, it is time-consuming for multi-camera systems to capture images with the calibrator in the ideal position. Hence, it is necessary for each camera to use as few images as possible to complete the calibration. Third, multi-camera systems need to be calibrated frequently because of the changing of camera position and lens. Besides, calibrator should be designed for ease of manufacturing while maintaining high machining precision to ensure the accuracy of calibration. As mentioned in section 3.2.2, degenerated cases lead the invalid solutions when all calibration points are coplanar. This also require that the calibrator must have the capability of providing non-coplanar three-dimensional calibration points.
To meet these requirements, we designed the calibrator as a cube. Each side of it is spray-painted with a different Apriltag label from the 36H11 family. Apriltag can provide five calibration points per plane, including four corners and a central point. Detecting more than two Apriltag patterns in each image can provide at least 10 points. These points also fulfill the condition of being non-coplanar, which avoids the degenerated cases mentioned in Section 3.2.2. Besides, the cube is easy to machine, with low cost but high machining accuracy. As for calibration points detection, Apriltag supports an open source algorithm, which is stable and easy to deploy.
Fig.4 illustrates the procedure for acquiring calibration data. Firstly, the cube is captured by all cameras in the shared field of view, with the center of the cube defined as the origin of the world coordinate system. The X-axis towards to number 1, the Y-axis towards to number 2 and the Z-axis towards to number 4. We define the set of images captured during this step is $Pack1$. Secondly, the cube is moved randomly to ensure that each camera captures an image containing at least two Apriltags, and the set of images captured in this phase is defined as $Pack2$. Until now, we have finished the calibration data collection. $Pack1$ is utilized for training the coarse extrinsic parameters, while $Pack2$ is used for calibrating coarse intrinsic parameters ${K_{cam}^{i}}$ of each camera.
Figure 5: Training sequential diagram: Our training process consists of three stages: 1) Camera Parameter Initialization Stage: The goal in this stage is to obtain coarse intrinsic and extrinsic parameters. Rendering training is not conducted in this stage. 2) Global Optimization Stage: Due to the coupling issue during joint optimization, we retain the Intrinsic Reprojection branch as a decoupling constraint. We also employ the progressive alignment method proposed in BARF to achieve extrinsic parameters with higher accuracy than those obtained in the first stage. 3) Fine-tuning Stage: In this stage, continuing to optimize extrinsic parameters without the progressive alignment constraint can lead to divergence. We freeze the extrinsic parameters and only adjust the intrinsic parameters.
3.3   Multi-Camera Extrinsic Parameters Regression
Once the coarse intrinsic parameters are obtained, the regression for extrinsic parameters is converted into a PNP problem which is Perspective-n-point. In the computation flow outlined in Section 3.2, the form of the loss function in Eq.(6) is extremely similar to that of the Bundle Adjustment (BA) method, which is widely employed to address Perspective-n-point (PnP) problems. This means that we can use the reprojection process to optimize the extrinsic parameters again.
3.3.1   World coordinate system
In Section 3.2, we collected two sets of auxiliary data, named $Pack1$ and $Pack2$, and the world coordinate system definitions in these two sets are entirely different. It’s important to note that the world coordinate system definition of the multi-camera acquisition system corresponds to that of $Pack1$. When using the $Pack2$ data to obtain camera intrinsic parameters, the matching extrinsic parameters can also be obtained. The origin of the world coordinate system for these extrinsic parameters remains defined at the center of the cube calibrator within their images. However, due to the random movement of the cube, the world coordinate systems for each image are not consistent with others. As a result, these extrinsic parameters are irrelevant to the multi-camera acquisition system and carry no practical significance. As illustrated in Fig.2, the $RT_{local}^{i}$ from Intrinsic reprojection loss branch is not reused in the subsequent optimization.
In order to obtain the accurate position of each camera in multi-camera acquisition system, the PnP optimization needs to adopt the Apritag data in $Pack1$ to regenerate extrinsic parameters.
3.3.2   PnP with bundle adjustment
The Perspective-n-point (PnP) problem is defined as follows: given a set of 3D points in world coordinate system, and knowing their corresponding coordinates in pixel coordinate system, solving the camera extrinsic parameters with the condition of intrinsic parameters provided. As a typical method to solve PnP, Bundle Adjustment (BA) can be described as to optimize camera poses and scene points by minimizing the difference between the projected 3D scene points in the images and their corresponding 2D image feature points. The loss function in BA is defined as:
Eq.(11) and Eq.(6) differ only in their coefficients, and the computation flow can be reused to regress the extrinsic parameters. The optimized coarse extrinsic parameters are denoted as $RT_{global}^{i}$.
During the training process, the two branches optimize simultaneously, sharing camera intrinsic parameters. Since their computation flows are almost the same, their convergence rates are quite similar, intrinsic and extrinsic reprojection losses can be directly added.
3.4   Neural Radiance Fields and Global Optimization
We unfold this section by explaining why NeRF is necessary for the global end-to-end network architecture first.
[…]
Then we discuss the coupling problem during joint training of intrinsic and extrinsic parameters. Following this, we validate the proposed method within the 2D image alignment space, which is similar to previous works such as BARF and L2G. We also explain the differences between jointly optimizing all camera parameters and individually optimizing extrinsic parameters in this part. Finally, we present the details of how we design the entire network and transition it from the 2D image alignment space to the 3D space.
3.4.1   Necessity of neural radiance fields
Neural radiance fields (NeRF) employ Multi-layer perceptron (MLP) to represent 3D scenes. The network processes 5D information, including position and direction, and generates pixel color and volume density from a given viewpoint.
An important precondition of NeRF is the accurate correspondence between image pixels and 3D space rays. On the one hand, we can use the camera intrinsic and extrinsic parameters to establish the above connection for 3D scene representation. On the other hand, we can also use this relationship as a constraint to optimize the camera parameters. Previous works, like SC-NeRF and BARF, have demonstrated the capability for jointly optimizing NeRF alongside camera parameters, which implies that it is feasible to implement an end-to-end joint optimization.
Our goal is to achieve a 3D representation of an object within the shared FOV of a multi-camera acquisition system, aligning with the inherent capabilities of NeRF. Additionally, it is an efficient and convenient method that can simultaneously acquire parameters for each camera during the scene representation training stage, with the ability to store and reuse them. This process seamlessly integrates calibration with scene reconstruction, offering significant benefits for multi-camera systems, particularly those experiencing frequent variations in intrinsic and extrinsic parameters.
In summary, the introduction of NeRF makes the end-to-end network structure design feasible, enabling simultaneous camera parameters regression and scene representation. These make NeRF necessary both in design and target.
3.4.2   Coupling in camera parameters
As mentioned in Eq.(1), our goal is to obtain the values of $K$ and $[R|T]$. In the NeRF method, rays are defined by their direction vectors rather than specific spatial point coordinates, which implies that we can disregard the scale variable $s$. Given pixel coordinates as inputs and $c=0$ provided in Eq.(4), Eq.(1) can be rewritten as follows:
${P_{i}}$ $={\left[{R\left|T\right.}\right]^{-1}}{K^{-1}}{p_{i}}$ (12)
[…]
Specifically, Eq.(12) can be modified as:
${P_{i}}$ $={R^{-1}}{K^{-1}}{p_{i}}-{R^{-1}}T$ (13)
$\Rightarrow{P_{i}}$ $=\tilde{R}{p_{i}}{\rm{+(}}\tilde{T}-{R^{-1}}T{\rm{)}}$
where $\tilde{R}$ and $\tilde{T}$ are as follows:
$\tilde{R}=\begin{bmatrix}r_{11}/f_{x}&r_{21}/f_{y}&r_{31}\\ r_{12}/f_{x}&r_{22}/f_{y}&r_{32}\\ r_{13}/f_{x}&r_{23}/f_{y}&r_{33}\end{bmatrix},$ (14)
$\tilde{T}=\begin{bmatrix}-u_{0}r_{11}/f_{x}&-v_{0}r_{21}/f_{y}&r_{31}\\ -u_{0}r_{12}/f_{x}&-v_{0}r_{22}/f_{y}&r_{32}\\ -u_{0}r_{13}/f_{x}&-v_{0}r_{23}/f_{y}&r_{33}\end{bmatrix}$
@@ 3.4.2   Coupling in camera parameters
$\tilde{T}-R^{-1}T$ $=\begin{bmatrix}\tilde{t}_{1}r_{11}+\tilde{t}_{2}r_{21}+\tilde{t}_{2}r_{31}\\ \tilde{t}_{1}r_{12}+\tilde{t}_{2}r_{22}+\tilde{t}_{2}r_{32}\\ \tilde{t}_{1}r_{13}+\tilde{t}_{2}r_{23}+\tilde{t}_{2}r_{33}\end{bmatrix}$ (15)

### v4

@@ 3.2.1   Coupling in camera parameters
$RGB$ and $\sigma$ denote the color and density of the sample points $\left\{{S_{n}}\right\}$ in 3D space. $\left\{{D_{n}}\right\}$ represents the view direction. $f_{sample}$ is the function that converts pixel coordinates into sampled points, and $pts_{n}$ is the pixel coordinate of pixel $n$. Specifically, $f_{sample}$ can be rewritten as follows:
$\left\{{{S_{n}}}\right\}=ray_{o}+ray_{d}\cdot\left\{{{s_{n}}}\right\}$ (3)
$\left\{{s_{n}}\right\}$ represents the sample points in standard space, which are related to the sampling distance. $ray_{o}$ denotes the origin and $ray_{d}$ represents the direction vector of the emitted ray. These items can be obtained as:
${ray_{o}}$ $=T\quad$ (4)
${ray_{d}}$ $=norm(Pts_{n}-T)\quad$
${Pts_{n}}$ $={[R|T]^{-1}}{K^{-1}}{pts_{n}}$
$Pts_{n}$ is 3D coordinate of $pts_{n}$ projected into 3D space. $norm$ represents the normalization operation. Moreover, ${Pts_{n}}$ in Eq.(4) can be modified as:
${Pts_{n}}$ $={[R|T]^{-1}}{K^{-1}}{pts_{n}}$ (5)
$\Rightarrow{Pts_{n}}$ $={R^{-1}}{K^{-1}}{pts_{n}}-{R^{-1}}T$
$\Rightarrow{Pts_{n}}$ $=\tilde{R}{pts_{n}}{\rm{+(}}\tilde{T}-{R^{-1}}T{\rm{)}}$
We define $[R|T]$ and $K$ as follows:
${[R|T]}$ $=\begin{bmatrix}r_{11}&r_{21}&r_{31}&t_{1}\\ r_{12}&r_{22}&r_{32}&t_{2}\\ r_{13}&r_{23}&r_{33}&t_{3}\end{bmatrix},$ (6)
$K$ $=\begin{bmatrix}f_{x}&0&u_{0}\\ 0&f_{y}&v_{0}\\ 0&0&1\end{bmatrix}$
where $\tilde{R}$ and $\tilde{T}$ in Eq.(5) are as follows:
$\tilde{R}=\begin{bmatrix}r_{11}/f_{x}&r_{12}/f_{y}&r_{13}\\ r_{21}/f_{x}&r_{22}/f_{y}&r_{23}\\ r_{31}/f_{x}&r_{32}/f_{y}&r_{33}\end{bmatrix},$ (7)
$\tilde{T}=\begin{bmatrix}-u_{0}r_{11}/f_{x}-v_{0}r_{12}/f_{y}\\ -u_{0}r_{21}/f_{x}-v_{0}r_{22}/f_{y}\\ -u_{0}r_{31}/f_{x}-v_{0}r_{32}/f_{y}\end{bmatrix}$
@@
$\tilde{T}-R^{-1}T$ $=\begin{bmatrix}\tilde{t}_{1}r_{11}+\tilde{t}_{2}r_{12}+\tilde{t}_{2}r_{13}\\ \tilde{t}_{1}r_{21}+\tilde{t}_{2}r_{22}+\tilde{t}_{2}r_{23}\\ \tilde{t}_{1}r_{31}+\tilde{t}_{2}r_{32}+\tilde{t}_{2}r_{33}\end{bmatrix}$ (8)